The perimeter of a swimming pool is 164meters. Determine the width of the pool if the length is 57 meters.
Answer by macston(5194)

.
P=perimeter=164m; L=length=57m; W=width
.
P=2(L+W)
P/2=L+W
P/2-L=W
(164m/2)-57m=W
82m-57m=W
25m=W
ANSWER: The width is 25 meters.